# Efficient Reasoning

### 05 Mar 2026 v1

@@ 2 Preliminaries
At the core of our approach is an explicit And/Or planning tree, which the agent constructs and maintains dynamically throughout the task. As new information is acquired, the agent continuously expands, revises, and prunes the planning tree to eliminate unpromising paths, enhancing computational efficiency while preventing unnecessary tree growth. Unlike conventional LLM agents, which lack the capability to autonomously construct and execute complex hierarchical plans, StructuredAgent separates planning responsibilities to improve both reliability and interpretability. Rather than delegating full tree construction and maintenance to the LLM, as in AgentOccam [Yang et al., 2025], StructuredAgent manages all aspects of tree construction directly, invoking the LLM only for targeted, well-defined operations including node expansion, repair, completion checking, and pruning. Constraining LLM queries to well-defined subtasks renders agent reasoning substantially more interpretable and robust. We provide detailed descriptions of these operations in Section˜3.3.
@@ 3.3 Tree Operations and Agent Modules
Example. We illustrate the modified greedy DFS on the task ‘‘Find a vegan chocolate brownie recipe with 4+ rating on allrecipes.com’’ (see Figure˜3). The agent begins with only the root node $1$ on the stack in Entering. In this example, Observation Summarizer and Global Tree Update are invoked after each Action node execution ($1.1$, $1.2.1$, $1.3$); Observation Summarizer updates the internal context while Global Tree Update makes no updates to the tree.
Steps 1–2 [Expand & Execute root and $1.1$] Node Expansion identifies node $1$ as an node and generates three ordered children: $1.1$ (type query), $1.2$ (select recipe), and $1.3$ (note recipe), all of which must succeed. Node $1$ is re-pushed in Exiting and its children are pushed in Entering in prescribed order. Node $1.1$ is then popped and identified as an Action node; the search query executes immediately and node $1.1$ is marked success and transitions through Exiting.
@@
Steps 5–6 [Execute $1.3$] Node $1.3$ is popped in Entering and executed as an Action node (noting the recipe). Node $1.3$ succeeds and transitions through Exiting. The root node $1$ is then popped in Exiting, and Node Completion Check confirms that all three children have succeeded. The stack empties and the agent terminates with success.
No node enters Failed in this trace, so repair and pruning are not triggered. Had node $1.2$ exhausted both Or alternatives without success, it would enter Failed, causing its unexecuted sibling $1.3$ to be pruned and failure to propagate upward to root node $1$, which would then invoke Node Repair as long as its revision budget is not exhausted.
@@ Appendix E Planning Tree
- [0] (AND) Research Pinterest to identify the top 3 women’s ethnic wear trends for Diwali. For each trend: 1) Specify the signature color currently trending. 2) Find one matching footwear style and one purse, both in the same color (same color as the trend) from the Pinterest Shop. 3) Provide the name of a potential brand or seller for each footwear and purse, and include a direct source or pin link if possible. Focus on vibrant, festive, and popular looks. Do not end your search early. Limit your review to only the first 2 scrolls of results. Do not click on external links. [success] - [0.1] (Atomic) Search for Diwali ethnic wear trends on Pinterest (type [401] [Diwali ethnic wear trends 2024] [1]) [success] - [0.2] (Atomic) Document 3 trending ethnic wear styles with their signature elements (note [Trending Diwali Ethnic Wear 2024: 1. Pink - Anupa Sari with embellished blouse from INDIASPOPUP.COM 2. Pista Green - Crush fabric sari with sequins embroidery border and yellow jacquard blouse from Etsy 3. Powder Peach - Stripe sari set with thread, pearl and sequin embroidery, featuring plunging neckline blouse from KYNAH]) [success] - [0.3] (AND) Navigate to Pinterest Shop and search for footwear in Pink, Pista Green, and Powder Peach shades from verified sellers specifically for Diwali looks [success] - [0.3.1] (Atomic) Navigate to Pinterest Shop section (click [208]) [success] - [0.3.2] (Atomic) type [401] [Pink ethnic juttis Diwali footwear] [1] (type [401] [Pink ethnic juttis Diwali footwear] [1]) [success] - [0.3.3] (Atomic) Return to previous page after completing pink footwear search (go_back) [success] - [0.3.4] (Atomic) type [401] [Pista Green ethnic juttis Diwali] [1] (type [401] [Pista Green ethnic juttis Diwali] [1]) [success] - [0.3.5] (Atomic) Return to previous page to continue accessory search (go_back) [success] - [0.3.6] (Atomic) type [401] [Powder Peach ethnic juttis Diwali] [1] (type [401] [Powder Peach ethnic juttis Diwali] [1]) [success] - [0.4] (OR) Search Pinterest Shop for matching ethnic clutches/purses in three colors by trying alternative search terms: 1) Search "Pink potli Diwali bag" or "Pink ethnic clutch" from verified sellers, 2) Search "Pista Green potli Diwali bag" or "Mint ethnic clutch" from verified sellers, 3) Search "Peach potli Diwali bag" or "Blush ethnic clutch" from verified sellers. Limit to first 2 scrolls of results. [success] - [0.4.1] (AND) Search for powder peach ethnic clutches in Shop section [success] - [0.4.1.1] (Atomic) Search for powder peach ethnic clutches in Shop section (type [401] [powder peach ethnic clutch Diwali] [1]) [success] - [0.4.1.2] (Atomic) Execute search query for powder peach ethnic clutches (type [401] [powder peach ethnic potli clutch bag Diwali] [1]) [success] - [0.4.1.4] (Atomic) note [Found matching powder peach potli bags from verified sellers: 1. 5 Elements Peach Brocade Pattern Potli Bag at Pernia’s Pop Up Shop 2. AMYRA Peach Poly Silk Pearl Embroidered Potli Bag 3. SWABHIMANN ACCESSORIES Peach Pure Silk Potli Bag] (note [Found matching powder peach potli bags from verified sellers: 1. 5 Elements Peach Brocade Pattern Potli Bag at Pernia’s Pop Up Shop 2. AMYRA Peach Poly Silk Pearl Embroidered Potli Bag 3. SWABHIMANN ACCESSORIES Peach Pure Silk Potli Bag]) [success] - [0.4.1.5] (Atomic) Modify search query to find matching powder peach potli bags (type [401] [powder peach potli bag Diwali] [1]) [success] - [0.4.1.6] (Atomic) click [208] to ensure Shop filter is active before searching for powder peach accessories (click [208] to ensure Shop filter is active before searching for powder peach accessories) [success] - [0.4.1.8] (Atomic) Modify search to find powder peach potli bags for Diwali (type [401] [powder peach potli bag Diwali] [1]) [success] - [0.5] (Atomic) Document powder peach accessory pairs with seller details (note [Powder Peach - Footwear: Coral Haze Juttis, Jutti Express formal wear, 5 Elements ethnic shoes; Purse: 5 Elements Brocade Potli with tassels, AMYRA Poly silk Pearl Embroidered Potli, Ri Ritu Kumar Net and Silk Zari Potli]) [success] - [0.6] (AND) Document pink and pista green accessory pairs with seller details and pin links, ensuring color coordination and authenticity of sellers [success] - [0.6.1] (AND) Search and document pink ethnic footwear with seller details [success] - [0.6.1.1] (Atomic) Search for pink ethnic footwear for Diwali (type [401] [pink ethnic juttis Diwali] [1]) [success] - [0.6.1.2] (Atomic) note [Found matching pink footwear options: 1. B Anu Designs - Baby Pink Punjabi Jutti with Mirror and Bead handwork 2. Miar Designs - Pink Leather Embroidered Juttis at Pernia’s Pop Up Shop 3. YASSIO - Pink Silk Floral Embroidered Juttis] (note [Found matching pink footwear options: 1. B Anu Designs - Baby Pink Punjabi Jutti with Mirror and Bead handwork 2. Miar Designs - Pink Leather Embroidered Juttis at Pernia’s Pop Up Shop 3. YASSIO - Pink Silk Floral Embroidered Juttis]) [success] - [0.6.1.3] (Atomic) Return to previous page to continue accessory search (go_back) [success] - [0.6.2] (Atomic) Document matching pink potli/clutch from search results (note [Found matching pink brocade potli bag from 5 Elements with tassels and beads embellishments, suitable for festive wear]) [success] - [0.6.3] (Atomic) Return to previous page after powder peach potli search (go_back) [success] - [0.6.4] (Atomic) Search for pista green ethnic footwear using specific search terms (type [401] [Pista green embroidered juttis Diwali] [1]) [success] - [0.6.5] (Atomic) Search for matching pista green clutch/potli (type [401] [Pista green brocade potli Diwali] [1]) [success] - [0.6.6] (Atomic) Document complete accessory details with seller information for pista green trend accessories: Footwear - House of Vian’s Crepe Embroidered Juttis and 5 Elements’ Leatherette Embellished Juttis; Purse - 5 Elements Green Brocade Embroidered Potli with tassels and dupion lining, and Tarini Nirula Green Brocade Stone Embellished Potli with antique gold frame (Document complete accessory details with seller information for pista green trend accessories: Footwear - House of Vian’s Crepe Embroidered Juttis and 5 Elements’ Leatherette Embellished Juttis; Purse - 5 Elements Green Brocade Embroidered Potli with tassels and dupion lining, and Tarini Nirula Green Brocade Stone Embellished Potli with antique gold frame) [success] - [0.6.7] (Atomic) Search for powder peach ethnic footwear (type [401] [powder peach ethnic juttis Diwali] [1]) [success] - [0.6.10] (AND) Search for pink ethnic footwear pins from verified sellers and save direct links [success] - [0.6.10.1] (Atomic) View details of pink leather embroidered juttis for Pink category (click [17999]) [success] - [0.6.10.2] (Atomic) note [Pink leather juttis by Miar Designs - Features: Leather base with coin, thread, pearl, sequins and beads embroidery, Available at Pernia’s Pop Up Shop, Price: 40,00, Formal occasion wear] (note [Pink leather juttis by Miar Designs - Features: Leather base with coin, thread, pearl, sequins and beads embroidery, Available at Pernia’s Pop Up Shop, Price: 40,00, Formal occasion wear]) [success] - [0.6.10.3] (Atomic) Return to previous page to continue purse search (go_back) [success] - [0.6.11] (Atomic) Search for matching pink potli/clutch pins (type [401] [pink ethnic potli clutch Diwali] [1]) [success] - [0.6.14] (Atomic) Verify and document color coordination between footwear-purse pairs (note [Color coordination verified: Pink category - Pink Beaded Mirror Work Potli and Pink Brocade Potli perfectly match the pink leather juttis from Miar Designs in shade and festive style]) [success] - [0.7] (Atomic) Document direct pin links for pink accessories from verified sellers (note [Found verified seller pin links: 1. Pink Brocade Tassels Embellished Potli by 5 Elements - Direct Pin ID 20730 2. Pink Brocade Ghungroo Embroidered Potli by 5 Elements - Direct Pin ID 20752 3. Pink Beaded Mirror Work Potli by Etsy verified seller - Direct Pin ID 21242]) [success] - [0.8] (AND) Review and document direct pin links for all three color categories ensuring each has verified footwear and purse pairs with seller details [success] - [0.8.1] (Atomic) Document completed findings for Pink category (note [Pink Category Complete - Footwear: Miar Designs, Purses: Pink Beaded Mirror Work Potli (Etsy), Pink Brocade Potli (5 Elements)]) [success] - [0.8.2] (Atomic) Navigate to Pinterest Shop section for ethnic potli clutches (click [20411]) [success] - [0.8.7] (Atomic) Search for powder peach potli options (type [401] [powder peach ethnic potli clutch Diwali] [1]) [success] - [0.8.8] (Atomic) Search for matching Pista Green potli options with verified sellers and clear product details (type [401] [pista green ethnic potli clutch Diwali] [1]) (Search for matching Pista Green potli options with verified sellers and clear product details (type [401] [pista green ethnic potli clutch Diwali] [1])) [success] - [0.8.9] (Atomic) Document verified seller information for Pista Green potlis (note [Verified Pista Green Potlis: 1. PLODE - Green Velvet Swarovski Triangle Potli with crystal embellishments and rhinestone strap 2. Velleso - Green Mirror Work Potli with crystal drops and silver threadwork (24x23cm) 3. The Pink Potli - Green Raw Silk Potli with bead work and tassel embellishments]) [success] - [0.8.10] (Atomic) Search for matching Powder Peach ethnic footwear options (type [401] [powder peach ethnic footwear Diwali] [1]) [success] - [0.8.11] (Atomic) Search for matching Pista Green ethnic footwear options with verified sellers and clear product details (type [401] [pista green ethnic footwear Diwali] [1]) [success] - [0.8.12] (Atomic) Document pin links and details for Pista Green category footwear and purse pairs (note [Found direct pin links for Pista Green category: 1) House of Vian Green Crepe Embroidered Juttis with crystals and gota patti from Pernia’s Pop Up Shop 2025, 2) 5 Elements Green Leatherette Embellished Juttis with beadwork from Pernia’s Pop Up Shop 2025]) [success]

### 06 Mar 2026 v2

@@ 2 Preliminaries
At the core of our approach is an explicit And/Or planning tree, which the agent constructs and maintains dynamically throughout the task. As new information is acquired, the agent continuously expands, revises, and prunes the planning tree to eliminate unpromising paths, enhancing computational efficiency while preventing unnecessary tree growth. Unlike conventional LLM agents, which lack the capability to autonomously construct and execute complex hierarchical plans, StructuredAgent separates planning responsibilities to improve both reliability and interpretability. Rather than delegating full tree construction and maintenance to the LLM, as in AgentOccam [Yang et al., 2025], StructuredAgent manages all aspects of tree construction directly, invoking the LLM only for targeted, well-defined operations including node expansion, repair, completion checking, and pruning. Constraining LLM queries to well-defined subtasks renders agent reasoning substantially more interpretable and robust. We provide detailed descriptions of these operations in Section˜$3.$
@@ 3.3 Tree Operations and Agent Modules
Example. We illustrate the modified greedy DFS on the task ‘‘Find a vegan chocolate brownie recipe with 4+ rating on allrecipes.com’’ (see Figure˜3). The agent begins with only the root node 11 on the stack in Entering. In this example, Observation Summarizer and Global Tree Update are invoked after each Action node execution ($1.1$, $1.2.1$, $1.3$); Observation Summarizer updates the internal context while Global Tree Update makes no updates to the tree.
Steps 1–2 [Expand & Execute root and $1.1$] Node Expansion identifies node 11 as an node and generates three ordered children: $1.1$ (type query), $1.2$ (select recipe), and $1.3$ (note recipe), all of which must succeed. Node 11 is re-pushed in Exiting and its children are pushed in Entering in prescribed order. Node $1.1$ is then popped and identified as an Action node; the search query executes immediately and node $1.1$ is marked success and transitions through Exiting.
@@
Steps 5–6 [Execute $1.3$] Node $1.3$ is popped in Entering and executed as an Action node (noting the recipe). Node $1.3$ succeeds and transitions through Exiting. The root node 11 is then popped in Exiting, and Node Completion Check confirms that all three children have succeeded. The stack empties and the agent terminates with success.
No node enters Failed in this trace, so repair and pruning are not triggered. Had node $1.2$ exhausted both Or alternatives without success, it would enter Failed, causing its unexecuted sibling $1.3$ to be pruned and failure to propagate upward to root node 11, which would then invoke Node Repair as long as its revision budget is not exhausted.
@@ Appendix E Planning Tree
- [0] (AND) Research Pinterest to identify the top 3 women’s ethnic wear trends for Diwali. For each trend: 1) Specify the signature color currently trending. 2) Find one matching footwear style and one purse, both in the same color (same color as the trend) from the Pinterest Shop. 3) Provide the name of a potential brand or seller for each footwear and purse, and include a direct source or pin link if possible. Focus on vibrant, festive, and popular looks. Do not end your search early. Limit your review to only the first 2 scrolls of results. Do not click on external links. [success] - [0.1] (Atomic) Search for Diwali ethnic wear trends on Pinterest (type [401] [Diwali ethnic wear trends 2024] [1]) [success] - [0.2] (Atomic) Document 3 trending ethnic wear styles with their signature elements (note [Trending Diwali Ethnic Wear 2024: 1. Pink - Anupa Sari with embellished blouse from INDIASPOPUP.COM 2. Pista Green - Crush fabric sari with sequins embroidery border and yellow jacquard blouse from Etsy 3. Powder Peach - Stripe sari set with thread, pearl and sequin embroidery, featuring plunging neckline blouse from KYNAH]) [success] - [0.3] (AND) Navigate to Pinterest Shop and search for footwear in Pink, Pista Green, and Powder Peach shades from verified sellers specifically for Diwali looks [success] - [0.3.1] (Atomic) Navigate to Pinterest Shop section (click [208]) [success] - [0.3.2] (Atomic) type [401] [Pink ethnic juttis Diwali footwear] [1] (type [401] [Pink ethnic juttis Diwali footwear] [1]) [success] - [0.3.3] (Atomic) Return to previous page after completing pink footwear search (go_back) [success] - [0.3.4] (Atomic) type [401] [Pista Green ethnic juttis Diwali] [1] (type [401] [Pista Green ethnic juttis Diwali] [1]) [success] - [0.3.5] (Atomic) Return to previous page to continue accessory search (go_back) [success] - [0.3.6] (Atomic) type [401] [Powder Peach ethnic juttis Diwali] [1] (type [401] [Powder Peach ethnic juttis Diwali] [1]) [success] - [0.4] (OR) Search Pinterest Shop for matching ethnic clutches/purses in three colors by trying alternative search terms: 1) Search "Pink potli Diwali bag" or "Pink ethnic clutch" from verified sellers, 2) Search "Pista Green potli Diwali bag" or "Mint ethnic clutch" from verified sellers, 3) Search "Peach potli Diwali bag" or "Blush ethnic clutch" from verified sellers. Limit to first 2 scrolls of results. [success] - [0.4.1] (AND) Search for powder peach ethnic clutches in Shop section [success] - [0.4.1.1] (Atomic) Search for powder peach ethnic clutches in Shop section (type [401] [powder peach ethnic clutch Diwali] [1]) [success] - [0.4.1.2] (Atomic) Execute search query for powder peach ethnic clutches (type [401] [powder peach ethnic potli clutch bag Diwali] [1]) [success] - [0.4.1.4] (Atomic) note [Found matching powder peach potli bags from verified sellers: 1. 5 Elements Peach Brocade Pattern Potli Bag at Pernia’s Pop Up Shop 2. AMYRA Peach Poly Silk Pearl Embroidered Potli Bag 3. SWABHIMANN ACCESSORIES Peach Pure Silk Potli Bag] (note [Found matching powder peach potli bags from verified sellers: 1. 5 Elements Peach Brocade Pattern Potli Bag at Pernia’s Pop Up Shop 2. AMYRA Peach Poly Silk Pearl Embroidered Potli Bag 3. SWABHIMANN ACCESSORIES Peach Pure Silk Potli Bag]) [success] - [0.4.1.5] (Atomic) Modify search query to find matching powder peach potli bags (type [401] [powder peach potli bag Diwali] [1]) [success] - [0.4.1.6] (Atomic) click [208] to ensure Shop filter is active before searching for powder peach accessories (click [208] to ensure Shop filter is active before searching for powder peach accessories) [success] - [0.4.1.8] (Atomic) Modify search to find powder peach potli bags for Diwali (type [401] [powder peach potli bag Diwali] [1]) [success] - [0.5] (Atomic) Document powder peach accessory pairs with seller details (note [Powder Peach - Footwear: Coral Haze Juttis, Jutti Express formal wear, 5 Elements ethnic shoes; Purse: 5 Elements Brocade Potli with tassels, AMYRA Poly silk Pearl Embroidered Potli, Ri Ritu Kumar Net and Silk Zari Potli]) [success] - [0.6] (AND) Document pink and pista green accessory pairs with seller details and pin links, ensuring color coordination and authenticity of sellers [success] - [0.6.1] (AND) Search and document pink ethnic footwear with seller details [success] - [0.6.1.1] (Atomic) Search for pink ethnic footwear for Diwali (type [401] [pink ethnic juttis Diwali] [1]) [success] - [0.6.1.2] (Atomic) note [Found matching pink footwear options: 1. B Anu Designs - Baby Pink Punjabi Jutti with Mirror and Bead handwork 2. Miar Designs - Pink Leather Embroidered Juttis at Pernia’s Pop Up Shop 3. YASSIO - Pink Silk Floral Embroidered Juttis] (note [Found matching pink footwear options: 1. B Anu Designs - Baby Pink Punjabi Jutti with Mirror and Bead handwork 2. Miar Designs - Pink Leather Embroidered Juttis at Pernia’s Pop Up Shop 3. YASSIO - Pink Silk Floral Embroidered Juttis]) [success] - [0.6.1.3] (Atomic) Return to previous page to continue accessory search (go_back) [success] - [0.6.2] (Atomic) Document matching pink potli/clutch from search results (note [Found matching pink brocade potli bag from 5 Elements with tassels and beads embellishments, suitable for festive wear]) [success] - [0.6.3] (Atomic) Return to previous page after powder peach potli search (go_back) [success] - [0.6.4] (Atomic) Search for pista green ethnic footwear using specific search terms (type [401] [Pista green embroidered juttis Diwali] [1]) [success] - [0.6.5] (Atomic) Search for matching pista green clutch/potli (type [401] [Pista green brocade potli Diwali] [1]) [success] - [0.6.6] (Atomic) Document complete accessory details with seller information for pista green trend accessories: Footwear - House of Vian’s Crepe Embroidered Juttis and 5 Elements’ Leatherette Embellished Juttis; Purse - 5 Elements Green Brocade Embroidered Potli with tassels and dupion lining, and Tarini Nirula Green Brocade Stone Embellished Potli with antique gold frame (Document complete accessory details with seller information for pista green trend accessories: Footwear - House of Vian’s Crepe Embroidered Juttis and 5 Elements’ Leatherette Embellished Juttis; Purse - 5 Elements Green Brocade Embroidered Potli with tassels and dupion lining, and Tarini Nirula Green Brocade Stone Embellished Potli with antique gold frame) [success] - [0.6.7] (Atomic) Search for powder peach ethnic footwear (type [401] [powder peach ethnic juttis Diwali] [1]) [success] - [0.6.10] (AND) Search for pink ethnic footwear pins from verified sellers and save direct links [success] - [0.6.10.1] (Atomic) View details of pink leather embroidered juttis for Pink category (click [17999]) [success] - [0.6.10.2] (Atomic) note [Pink leather juttis by Miar Designs - Features: Leather base with coin, thread, pearl, sequins and beads embroidery, Available at Pernia’s Pop Up Shop, Price: 40,$0$, Formal occasion wear] (note [Pink leather juttis by Miar Designs - Features: Leather base with coin, thread, pearl, sequins and beads embroidery, Available at Pernia’s Pop Up Shop, Price: 40,$0$, Formal occasion wear]) [success] - [0.6.10.3] (Atomic) Return to previous page to continue purse search (go_back) [success] - [0.6.11] (Atomic) Search for matching pink potli/clutch pins (type [401] [pink ethnic potli clutch Diwali] [1]) [success] - [0.6.14] (Atomic) Verify and document color coordination between footwear-purse pairs (note [Color coordination verified: Pink category - Pink Beaded Mirror Work Potli and Pink Brocade Potli perfectly match the pink leather juttis from Miar Designs in shade and festive style]) [success] - [0.7] (Atomic) Document direct pin links for pink accessories from verified sellers (note [Found verified seller pin links: 1. Pink Brocade Tassels Embellished Potli by 5 Elements - Direct Pin ID 20730 2. Pink Brocade Ghungroo Embroidered Potli by 5 Elements - Direct Pin ID 20752 3. Pink Beaded Mirror Work Potli by Etsy verified seller - Direct Pin ID 21242]) [success] - [0.8] (AND) Review and document direct pin links for all three color categories ensuring each has verified footwear and purse pairs with seller details [success] - [0.8.1] (Atomic) Document completed findings for Pink category (note [Pink Category Complete - Footwear: Miar Designs, Purses: Pink Beaded Mirror Work Potli (Etsy), Pink Brocade Potli (5 Elements)]) [success] - [0.8.2] (Atomic) Navigate to Pinterest Shop section for ethnic potli clutches (click [20411]) [success] - [0.8.7] (Atomic) Search for powder peach potli options (type [401] [powder peach ethnic potli clutch Diwali] [1]) [success] - [0.8.8] (Atomic) Search for matching Pista Green potli options with verified sellers and clear product details (type [401] [pista green ethnic potli clutch Diwali] [1]) (Search for matching Pista Green potli options with verified sellers and clear product details (type [401] [pista green ethnic potli clutch Diwali] [1])) [success] - [0.8.9] (Atomic) Document verified seller information for Pista Green potlis (note [Verified Pista Green Potlis: 1. PLODE - Green Velvet Swarovski Triangle Potli with crystal embellishments and rhinestone strap 2. Velleso - Green Mirror Work Potli with crystal drops and silver threadwork (24x23cm) 3. The Pink Potli - Green Raw Silk Potli with bead work and tassel embellishments]) [success] - [0.8.10] (Atomic) Search for matching Powder Peach ethnic footwear options (type [401] [powder peach ethnic footwear Diwali] [1]) [success] - [0.8.11] (Atomic) Search for matching Pista Green ethnic footwear options with verified sellers and clear product details (type [401] [pista green ethnic footwear Diwali] [1]) [success] - [0.8.12] (Atomic) Document pin links and details for Pista Green category footwear and purse pairs (note [Found direct pin links for Pista Green category: 1) House of Vian Green Crepe Embroidered Juttis with crystals and gota patti from Pernia’s Pop Up Shop 2025, 2) 5 Elements Green Leatherette Embellished Juttis with beadwork from Pernia’s Pop Up Shop 2025]) [success]